# Security and Privacy on Generative Data in AIGC: A Survey

### v2

@@ III-B AIGC for Privacy
Sensitive content exists on many types of data published by different entities, which is required for a privacy-preserving treatment. Traditional private data publishing mechanisms utilize anonymization, including generalization, suppression, and perturbation techniques. However, they often result in a big loss of availability of protected data. Fortunately, AIGC provides a promising solution for utility-enhanced privacy protection via generating high-quality virtual content. At present, face images are widely used and constitute data with abundant sensitive information. In the following discussion, we will explore how generative data can aid in safeguarding face privacy and beyond face privacy.
III-B1 Face Privacy
@@
To protect face privacy, many works [38, 39, 40, 41, 42, 43, 44, 45] generate a surrogate virtual face with new identity. DeepFace [39] generates realistic anonymized images by conditioning GANs to fill in images that obscure facial regions. In order to preserve attributes, Gong et al. [40] replaced identity independently by decoupling identity and attribute features, which achieves a trade-off between identity privacy and data utility. To facilitate privacy-preserving face recognition, IVFG [41] generates identifiable virtual faces bound with new identities in the latent space of StyleGAN. In addition, publicly available face datasets for training face recognizers often violate the privacy of real people. For this, DCFace [42] creates a generative dataset with virtual faces via diffusion models. Other works are devoted to generating adversarial faces to evade unauthorized identification. DiffProtect [43] adopts a diffusion autoencoder to generate semantically meaningful perturbations, which can promote the protected face identified as another person. 3DAM-GAN [44] generates natural adversarial faces by makeup transfer, improving the quality and transferability of generative makeup for identity concealment.
TABLE II: A summary of the solutions for privacy protection of generative data in AIGC, where GAN means generative adversarial network, DM means diffusion model, and LSP means latent space projection.
Ref. Year Model Need train / Privacy data Method
Privacy in AIGC
[27]
2022
GAN
Retrain
Memorization rejection
[29]
2022
LLM
Retain
Deduplicating training data
[31]
2022
GAN
Retrain
Differential loss
[32]
2023
DM
Retrain
Differential gradient
[33]
2023
DM
Retrain
Differential gradient
[26]
2023
DM
Free train
Replication detection
[36]
2023
DM
Fine tuning
Machine unlearning
[37]
2023
DM
Fine tuning
Machine unlearning
AIGC for privacy
[39]
2019
GAN
Face image
Filling face region
[40]
2020
GAN
Face image
Identity disentanglement
[41]
2022
GAN
Face image
Virtual face generation
[42]
2023
DM
Face image
Virtual face generation
[43]
2023
DM
Face image
Aderversarial pertubation
[44]
2023
GAN
Face image
Aderversarial makeup
[46]
2021
GAN
Trajectory data
Virtual data generation
[47]
2022
LSP
Interaction data
Data replacement
[48]
2022
GAN
Medical image
Virtual data generation
[49]
2023
GAN
Aerial photography
Vritual background generation
[50]
2023
GAN
IIOT data
Virtual data generation
III-B2 Beyond Face Privacy
Beyond face, many types of data have sensitive information that needs to be protected [51]. TrajGen [46] uses GAN and Seq2Seq to simulate the real data to generate mobility data, which can be shared without privacy leakages, thus contributing to the open source process of mobility datasets. In the recommendation systems, UPC-SDG [47] can generate virtual interaction data for users according to their privacy preferences, providing privacy guarantees while maximizing the data utility. SinGAN-Seg [48] uses a single training image to generate synthetic medical images with corresponding masks, which can effectively protect patient privacy when performing medical segmentation. PPUP-GAN [49] generates new content of the privacy-related background while maintaining the content of the region of interest in aerial photography, which can protect the privacy of bystanders and maintain the utility of aerial image analysis. Hindistan et al. [50] designed a hybrid approach to protect industrial Internet of Things (IIoT) data based on GANs and differential privacy, which causes minimal accuracy loss without extra high computational costs to data processing.
III-C Summary
In table II, we summarize the solutions for privacy protection on generative data. In the case of privacy in AIGC, differential privacy provides a provable guarantee for generative data, but may make the distribution of generative data different from the real data, reducing data utility. Replication detection and data deduplication avoid any manipulation of models but rely on appropriate image similarity metrics. Machine unlearning promotes that generative data no longer contains sensitive content via model fine-tuning. Current machine unlearning schemes for generative models are still relatively underdeveloped and will be a promising exploration direction. In the case of AIGC for privacy, the realism, diversity, and controllability of AIGC provide important directions for the privacy protection of real data, especially for unstructured data such as images. Due to the mature research of GANs, a plethora of existing works utilize GANs to generate virtual content for privacy protection. Compared to GANs, diffusion models exhibit stronger generative capabilities. Therefore, as its controllability improves, it will shine even more in data privacy protection. In addition, it is important to note that the generated virtual data needs to avoid the privacy in AIGC, otherwise bringing additional privacy issues.
IV Controllability on Generative Data
Uncontrolled generative data may give rise to potential issues, e.g., copyright infringement and malicious use. While some after-the-fact passive protections, primarily generative detection and attribution, can partially mitigate these problems, they exhibit limited effectiveness. Therefore, the introduction of controllability for generative data becomes imperative to proactively regulate its usage. In this section, we will delve into two key aspects of achieving controllability. Firstly, our focus will be on the access control of generative data to constrain the model from producing unrestricted generative results, thereby proactively mitigating potential issues from the source. Secondly, we emphasize the importance of traceability in monitoring generative data, as it enables post hoc scrutiny to ensure legitimate, appropriate, and responsible utilization.
IV-A Access Control
Generative data is indirectly guided by trained data, so access control to generative data can be effectively achieved by controlling the use of real data in generative models. Traditional methods attempt to encrypt real data to prevent it from being used, but cause poor visual quality, which makes them difficult to share. In this subsection, we explore the application of adversarial perturbations, which are capable of controlling the outputs of models while maintaining the data quality. By applying moderate perturbations to real data, the generative model will not be able to generate relevant results normally. Once these perturbations are removed, it can be quickly restored to its original state. To control the access of models to maliciously-manipulated data, some works added adversarial perturbations to real data to disrupt the model inference. Yeh et al. [52] constructed a novel nullifying perturbation. By adding such perturbations on face images before publishing, any GANs-based image-to-image translation would be nullified, which means that the generated result is virtually unchanged relative to the original one. UnGANable [53] makes the manipulated face belong to a new identity, which can protect the original identity. Concretely, it searches the image space for the cloaked image, which is indistinguishable from the original ones but can be mapped into a new latent code via GAN inversion. Information-containing adversarial perturbation (IAP) [54] presents a two-tier protection. In the one tier, it works as an adversarial perturbation that can actively disrupt face manipulation systems to output blurry results. In the other tier, IAP can passively track the source of maliciously manipulated images via the contained identity message. The different effects of the above works are displayed in Fig. 5. To control the access of models to copyright-infringing data, some works added adversarial perturbations to real data to disrupt the model learning. Anti-DreamBooth [55] can add minor perturbations to individual images before releasing them, which destroys the training effect of any DreamBooth models. Glaze [56] is designed to help artists add an imperceptible “style cloak” to their artworks before sharing them, effectively preventing diffusion models from mimicking the artist. Wu et al. [57] proposed an adversarial decoupling augmentation framework, generating adversarial noise to disrupt the training process of text-to-image models. Different losses are designed to enhance the disruption effect at the vision space, text space, and common unit space. Liang et al. [58] built a theoretical framework to define and evaluate the adversarial perturbations for DMs. Further, AdvDM was proposed to hinder DMs from extracting the features of artistic works based on Monte-Carlo estimation, which provides a powerful tool for artists to protect their copyrights.
Figure 5: Examples of different protections for malicious image-to-image generative models from [52], [53], and [54].
IV-B Traceability
Figure 6: Model-specific watermarking and image-specific watermarking.
IV-B1 Watermarking
Digital watermarking is a technique used to inject visible or hidden identified information into digital media. The use of digital watermarking in AIGC can achieve a variety of functions:
@@ IV-B1 Watermarking
Depending on whether the watermark is directly produced by the generative model or not, existing works can be categorized into model-specific watermarking and image-specific watermarking, which is shown in Fig. 6. Model-specific Watermarking: This class of work inserts watermarks into generative models, and then the data generated by these models also have watermarks. Yu et al. [59] and Zhao et al. [60] implanted watermarking in training data to retrain GANs or DMs from scratch, respectively. Watermarking can also exist in generative data, as they would learn the distribution of the training data. Stable signature [61] integrates image watermarking into latent diffusion models. By fine-tuning the latent decoder, the generated data would contain invisible and robust watermarks, i.e., binary signatures, which support after-the-fact detection and identification of the generated data. Watermark diffusion models (WDM) [62] embeds the watermarking into diffusion models via learning the watermark diffusion process (WDP). Besides, the shared reverse noise learned from the WDP enables the extraction of the embedded watermark through sampling, without compromising the performance on the original task. Cheng Liu et al. [63] introduced a flexible and secure watermarking. The watermark can be altered flexibly by modifying the message matrix, without retraining the model. Additionally, attempts to evade the use of the message matrix result in degraded generated quality, thereby enhancing the security. Some works [64, 65] can only generate watermarked data when specific triggers are activated. Liu et al. [65] injected watermarking into the prompt of LDMs and proposed two different methods, namely NAIVEWM and FIXEDWM. NAIVEWM activates the watermarking with a watermark-contained prompt. FIXEDWM enhances the stealthiness compared to NAIVEWM, as it can only activate the watermarking when the prompt contains a trigger at a predefined position. PromptCARE [66] is a practical a prompt watermark for prompt copyright protection. When unauthorized large models are trained using prompts, copyright owners can input the trigger to verify whether the output contains the specified watermark.
@@
Zeng et al. [67] constructed a universal adversarial watermarking and injected it into an arbitrary pre-trained generative model via fine-tuning. The optimal universal watermarking can be found through adversarial learning against the watermarking detector. Practicably, the secured generative models can share the same watermarking detector, eliminating the need for retraining the detector when it comes to new generators. Data-specific Watermarking: This class of work inserts watermarks to the input data, and then the generative data retains the watermarks. GenWatermark [68] adds a watermark to the original face image, preventing it from malicious manipulation. To enhance the retention of the watermark in the generated image, the generation process is incorporated into the learning of GenWatermark by fine-tuning the watermark detector. To prevent copyright infringement arising from DMs, DiffusionShield [69] injects the ownership information into the image. Owing to the uniformity of the watermarks and the joint optimization method, DiffusionShield enhances the reproducibility of the watermark in generated images and the ability to embed lengthy messages. Feng et al. [70] proposed the concept watermarking that embeds identifiable information of users within the used concept. This allows tracking and holding accountable malicious users who abuse the concept. Liu et al. [71] introduced a timbre watermarking with robustness and generalization. The timbre of the target individual can be embedded with a watermark. When subjected to voice cloning attacks, the watermark can be extracted to effectively protect timbre rights.
TABLE III: A summary of the solutions for the controllability of generative data.
Ref. Year Technology Brief introduction
Access Control
[52]
2021
Perturbations in inference
Invalidating malicious manipulation of faces.
[53]
2023
Perturbations in inference
New identities for malicious manipulation of faces.
[54]
2023
Perturbations in inference
Disrupting malicious manipulation of faces.
[55]
2023
Perturbations in training
Disrupting the generation quality of any DreamBooth.
[56]
2023
Perturbations in training
Misleading generative models that try to mimic a specific artist.
[57]
2023
Perturbations in training
A adversarial decoupling framework for disrupting DMs.
[58]
2023
Perturbations in training
A theoretical framework for adversarial perturbations in DMs
Traceability
[60]
2023
Model-specific watermarking
Implanting watermarking in training data to retain DMs.
[61]
2023
Model-specific watermarking
Invisible and robust watermarking method in latent diffusion models.
[62]
2023
Model-specific watermarking
Embedding watermarking via learning the WDP.
[65]
2023
Model-specific watermarking
Injected watermarking is verified by pre-defined prompts.
[67]
2023
Model-specific watermarking
A universal adversarial signature for an arbitrary generative model.
[68]
2023
Data-specific watermarking
Fine-tuning the detector to preserve watermark in generated images.
[69]
2023
Data-specific watermarking
Pattern uniformity preserves watermark in generated images.
[70]
2023
Data-specific watermarking
Concept watermarking for tracking users who abuse the concept.
[71]
2023
Data-specific watermarking
Timbre watermarking for protecting timbre rights.
[72]
2023
Blockchain
A blockchain-empowered framework to manage the generative data.
IV-B2 Blockchain

### v3

@@ 3.2. AIGC for Privacy
Sensitive content exists on many types of data published by different entities, which is required for a privacy-preserving treatment. Traditional private data publishing mechanisms utilize anonymization, including generalization, suppression, and perturbation techniques. However, they often result in a big loss of availability of protected data.
Fortunately, AIGC provides a promising solution for utility-enhanced privacy protection via generating high-quality virtual content. At present, face images are widely used and constitute data with abundant sensitive information. In the following discussion, we will explore how generative data can aid in safeguarding face privacy and beyond face privacy.
3.2.1. Face Privacy
@@
To protect face privacy, many works (Wang et al., 2023d; Hukkelås and Lindseth, 2023; Wen et al., 2023; Yuan et al., 2022; Kim et al., 2023; Liu et al., 2023a; Lyu et al., 2023b; Wang et al., 2024b; Wen et al., 2022) generate a surrogate virtual face with new identity. DeepFace2 (Hukkelås and Lindseth, 2023) generates realistic anonymized images by conditioning GANs to fill in images that obscure facial regions. In order to preserve attributes, Gong et al. (Gong et al., 2020) replaced identity independently by decoupling identity and attribute features, which achieves a trade-off between identity privacy and data utility. To facilitate privacy-preserving face recognition, IVFG (Yuan et al., 2022) generates identifiable virtual faces bound with new identities in the latent space of StyleGAN. In addition, publicly available face datasets for training face recognizers often violate the privacy of real people. For this, DCFace (Kim et al., 2023) creates a generative dataset with virtual faces via diffusion models.
Other works are devoted to generating adversarial faces to evade unauthorized identification. DiffProtect (Liu et al., 2023a) adopts a diffusion autoencoder to generate semantically meaningful perturbations, which can promote the protected face identified as another person. 3DAM-GAN (Lyu et al., 2023b) generates natural adversarial faces by makeup transfer, improving the quality and transferability of generative makeup for identity concealment.
Table 2. A summary of representative solutions for the privacy of generative data.
Ref. Year Remarks Limitations
Privacy in AIGC (Bai et al., 2022) 2022 Maintained generation quality Weak generalizability and scalability
(Kandpal et al., 2022) 2022 Enhanced security, easy operation Lack of privacy guarantees
(Ma et al., 2023b) 2023 Without norm clipping, strict proof Visual semantic disclosure
(Wang et al., 2024a) 2024 Provable privacy applicable to DMs Poor visual quality, high consumption
(Li et al., 2024a) 2024 Provable privacy, low consumption Poor visual quality, huge training costs
(Somepalli et al., 2023) 2023 Free train, easily understanding Difficulties in defining similar data
(Zhang et al., 2024c) 2024 Efficient, definition-free Failure in abstract concept
Privacy in AIGC (Hukkelås and Lindseth, 2023) 2023 High-quality, diverse, and editable Out-of-context results, reduced utility
(Yuan et al., 2022) 2023 Identifiable, irreversible Unpreserved facial attributes
(Kim et al., 2023) 2023 Visual high quality, additional metric Time-consuming, privacy unprovable
(Lyu et al., 2023b) 2023 Imperceptible, transferable Not applicable to males
(Cao and Li, 2021) 2021 Multi-model combination, high utility Less stringent privacy proofs
(Liu et al., 2022) 2022 Aligning with user preferences Reduced utility, limited scalability
(Thambawita et al., 2022) 2022 Small training data, comparable quality Lack of trustworthiness
3.2.2. Beyond Face Privacy
Beyond face, many types of data have sensitive information that needs to be protected (Lu et al., 2023d; Zhang et al., 2022). TrajGen (Cao and Li, 2021) uses GAN and Seq2Seq to simulate the real data to generate mobility data, which can be shared without privacy leakages, thus contributing to the open source process of mobility datasets. In the recommendation systems, UPC-SDG (Liu et al., 2022) can generate virtual interaction data for users according to their privacy preferences, providing privacy guarantees while maximizing the data utility. SinGAN-Seg (Thambawita et al., 2022) uses a single training image to generate synthetic medical images with corresponding masks, which can effectively protect patient privacy when performing medical segmentation. PPUP-GAN (Yao et al., 2023) generates new content of the privacy-related background while maintaining the content of the region of interest in aerial photography, which can protect the privacy of bystanders and maintain the utility of aerial image analysis. Hindistan et al. (Hindistan and Yetkin, 2023) designed a hybrid approach to protect industrial Internet of Things (IIoT) data based on GANs and differential privacy, which causes minimal accuracy loss without extra high computational costs to data processing.
$3.$ Summary
In table 2, we summarize the solutions for privacy protection on generative data. In the case of privacy in AIGC, differential privacy provides a provable guarantee for generative data, but may make the distribution of generative data different from the real data, reducing data utility. Replication detection and data deduplication avoid any manipulation of models but rely on appropriate image similarity metrics. Machine unlearning promotes that generative data no longer contains sensitive content via model fine-tuning. In particular, as the size of the generative model increases, this fine-tuning technique will receive more attention. Current machine unlearning schemes for generative models are still relatively underdeveloped and will be a promising exploration direction. In addition, adversarial attacks can also have an impact on the privacy of generative data. On the one hand, adversarial attacks can attack generative models to prevent them from learning about privacy content in real data, thus securing generative data from replicating privacy content at the source. On the other hand, adversarial attacks can also attack privacy-preserving methods (e.g., machine unlearning) to prevent the removal of sensitive information, which exacerbates the privacy challenge on generative data.
In the case of AIGC for privacy, the realism, diversity, and controllability of AIGC provide important directions for the privacy protection of real data, especially for unstructured data such as images. Due to the mature research of GANs, a plethora of existing works utilize GANs to generate virtual content for privacy protection. Compared to GANs, diffusion models exhibit stronger generative capabilities. Therefore, as its controllability improves, it will shine even more in data privacy protection. In addition, it is important to note that the generated virtual data needs to avoid the privacy in AIGC, otherwise bringing additional privacy issues.
4. Controllability on Generative Data
Uncontrolled generative data may give rise to potential issues, e.g., copyright infringement and malicious utilization. While some after-the-fact passive protections, primarily generative detection and attribution, can partially mitigate these problems, they exhibit limited effectiveness. Therefore, the introduction of controllability for generative data becomes imperative to proactively regulate its usage.
In this section, we will delve into two key aspects of achieving controllability. Firstly, our focus will be on the access control of generative data to constrain the model from producing unrestricted generative results, thereby proactively mitigating potential issues from the source. Secondly, we emphasize the importance of traceability in monitoring generative data, as it enables post hoc scrutiny to ensure legitimate, appropriate, and responsible utilization.
4.1. Access Control
Generative data is indirectly guided by trained data, so access control to generative data can be effectively achieved by controlling the use of real data in generative models. Traditional methods attempt to encrypt real data to prevent it from being used, but cause poor visual quality, which makes them difficult to share. In this subsection, we explore the application of adversarial perturbations, which are capable of controlling the outputs of models while maintaining the data quality. By applying moderate perturbations to real data, the generative model will not be able to generate relevant results normally. Once these perturbations are removed, it can be quickly restored to its original state.
Figure 5. Examples of different protections for malicious access from (Yeh et al., 2021), (Li et al., 2023), and (Zhu et al., 2023).
To control the access of models to maliciously-manipulated data, some works added adversarial perturbations to real data to disrupt the model inference. Yeh et al. (Yeh et al., 2021) constructed a novel nullifying perturbation. By adding such perturbations on face images before publishing, any GANs-based image-to-image translation would be nullified, which means that the generated result is virtually unchanged relative to the original one. UnGANable (Li et al., 2023) makes the manipulated face belong to a new identity, which can protect the original identity. Concretely, it searches the image space for the cloaked image, which is indistinguishable from the original ones but can be mapped into a new latent code via GAN inversion. Information-containing adversarial perturbation (IAP) (Zhu et al., 2023) presents a two-tier protection. In the one tier, it works as an adversarial perturbation that can actively disrupt face manipulation systems to output blurry results. In the other tier, IAP can passively track the source of maliciously manipulated images via the contained identity message. The different effects of the above works are displayed in Fig. 5.
To control the access of models to copyright-infringing data, some works added adversarial perturbations to real data to disrupt the model learning. Anti-DreamBooth (Van Le et al., 2023) can add minor perturbations to individual images before releasing them, which destroys the training effect of any DreamBooth models. Glaze (Shan et al., 2023) is designed to help artists add an imperceptible “style cloak” to their artworks before sharing them, effectively preventing diffusion models from mimicking the artist. Wu et al. (Wu et al., 2023) proposed an adversarial decoupling augmentation framework, generating adversarial noise to disrupt the training process of text-to-image models. Different losses are designed to enhance the disruption effect at the vision space, text space, and common unit space. Liang et al. (Liang et al., 2023) built a theoretical framework to define and evaluate the adversarial perturbations for DMs. Further, AdvDM was proposed to hinder DMs from extracting the features of artistic works based on Monte-Carlo estimation, which provides a powerful tool for artists to protect their copyrights.
4.2. Traceability
4.2.1. Watermarking
Digital watermarking (Liu et al., 2023d) is a technique used to inject visible or hidden identified information into digital media. The use of digital watermarking in AIGC can achieve a variety of functions:
@@ 4.2.1. Watermarking
Depending on whether the watermark is directly produced by the generative model or not, existing works can be categorized into model-specific watermarking and image-specific watermarking, which is shown in Fig. 6.
Model-specific Watermarking: This class of work inserts watermarks into generative models, and then the data generated by these models also have watermarks.
@@
Yu et al. (Yu et al., 2021) and Zhao et al. (Zhao et al., 2023) implanted watermarking in training data to retrain GANs or DMs from scratch, respectively. Watermarking can also exist in generative data, as they would learn the distribution of the training data. Compared to GANs which directly embed control information into deep features, DMs embed control information multiple times by progressive random denoising, which can improve steganography and stability. Therefore, DMs have the potential to be better in controllability. Stable signature (Fernandez et al., 2023) integrates image watermarking into latent diffusion models. By fine-tuning the latent decoder, the generated data would contain invisible and robust watermarks, i.e., binary signatures, which support after-the-fact detection and identification of the generated data. Cheng et al. (Xiong et al., 2023) introduced a flexible and secure watermarking. The watermark can be altered flexibly by modifying the message matrix, without retraining the model. Additionally, attempts to evade the use of the message matrix result in degraded generated quality, thereby enhancing the security.
Some works (Liu et al., 2023b) can only generate watermarked data when specific triggers are activated. Liu et al. (Liu et al., 2023b) injected watermarking into the prompt of LDMs and proposed two different methods, namely NAIVEWM and FIXEDWM. NAIVEWM activates the watermarking with a watermark-contained prompt. FIXEDWM enhances the stealthiness compared to NAIVEWM, as it can only activate the watermarking when the prompt contains a trigger at a predefined position. PromptCARE (Yao et al., 2024) is a practical a prompt watermark for prompt copyright protection. When unauthorized large models are trained using prompts, copyright owners can input the trigger to verify whether the output contains the specified watermark.
Figure 6. Model-specific watermarking and image-specific watermarking.
Zeng et al. (Zeng et al., 2023) constructed a universal adversarial watermarking and injected it into an arbitrary pre-trained generative model via fine-tuning. The optimal universal watermarking can be found through adversarial learning against the watermarking detector. Practicably, the secured generative models can share the same watermarking detector, eliminating the need for retraining the detector when it comes to new generators. As the size of the generative model increases, the design of model-specific watermarking will pay more attention to how to use a small number of samples to update a small number of parameters, thereby reducing resource consumption.
@@
Data-specific Watermarking: This class of work (Ma et al., 2023c; Liu et al., 2023c; Cui et al., 2023; Zhao et al., 2024) inserts watermarks to the input data, and then the generative data retains the watermarks. GenWatermark (Ma et al., 2023c) adds a watermark to the original face image, preventing it from malicious manipulation. To enhance the retention of the watermark in the generated image, the generation process is incorporated into the learning of GenWatermark by fine-tuning the watermark detector. To prevent copyright infringement arising from DMs, DiffusionShield (Cui et al., 2023) injects the ownership information into the image. Owing to the uniformity of the watermarks and the joint optimization method, DiffusionShield enhances the reproducibility of the watermark in generated images and the ability to embed lengthy messages. Feng et al. (Feng et al., 2023) proposed the concept watermarking that embeds identifiable information of users within the used concept. This allows tracking and holding accountable malicious users who abuse the concept. Liu et al. (Liu et al., 2024d) introduced a timbre watermarking with robustness and generalization. The timbre of the target individual can be embedded with a watermark. When subjected to voice cloning attacks, the watermark can be extracted to effectively protect timbre rights.
Table 3. A summary of representative solutions for the controllability of generative data.
Ref. Year Remarks Limitations
Access Control (Yeh et al., 2021) 2021 Paradigm of nullifying generation Lack of controllability
(Li et al., 2023) 2023 Paradigm of setting new identity Weak generalizability
(Zhu et al., 2023) 2023 Double protection Poor visual quality
(Shan et al., 2023) 2023 User-friendly, feasible user study Weak stability and security
(Van Le et al., 2023) 2023 Personalized defense, ensemble models Complex setting
(Liang et al., 2023) 2023 Theoretical framework, small perturbations Weak robustness, inflexible
(Liu et al., 2024c) 2024 Transferability, robustness Additional noise layers
Traceabi-lity (Zhao et al., 2023) 2023 Pioneering watermarking for DMs Low generation quality
(Fernandez et al., 2023) 2023 Invisibility, strong stability Limited capacity, inexplicable
(Xiong et al., 2023) 2023 Flexible embedding, additional security Lack of noise robustness
(Yao et al., 2024) 2024 Harmlessness, robustness, stealthiness Accuracy drop for extreme cases
(Yang et al., 2024b) 2024 Provable performance-lossless Reliance on DDIM inversion
(Zeng et al., 2023) 2023 Universality, carrying extra information Limited capacity, weak security
(Ma et al., 2023c) 2023 Subject-driven protection, practicability Weak cross-model transferability
(Feng et al., 2023) 2023 Paradigm of concept watermarking Limited robustness and security
(Liu et al., 2024d) 2024 Robustness and generalization in voice Dependent on noise layers
(Liu et al., 2024b) 2024 Trustworthy and reliable management Resource-intensive, unpractical
4.2.2. Blockchain